In 1994 the population of Hannaford, MO increased by 20%. In 1995 it increased
16% from 1994. The population is now 22,272. Was there a larger increase in
people in 1994 or 1995, and by how much?

Answer by Theo(3458)

if now is 1995, then this problem can be solved as shown.
if not, we would need to know the increase from 1995 to now (2009).
since that information wasn't part of the problem, I assumed now is 1995.

1994 = 1993 * 1.2
1995 = 1994 * 1.16
1995 = 22272
1994 = 22272/1.16 = 19200
1993 = 19200/1.2 = 16000

you get:

1993 = 16000
1994 = 16000 * 1.2 = 19200
1995 = 19200 * 1.16 = 22272

increase 1994 over 1993 = 19200 - 16000 = 3200

increase 1995 over 1994 = 22272 - 19200 = 3072

there was a larger increase in number of people in 1994 over 1995.

the increase was 3200-3072 = 128 more people increase in 1994 than in 1995.